OpenStudy (anonymous):

Peter played his favorite video game for 10 hours last week. Today Peter's parents restricted him to 5 hours each week for the next 8 weeks. Create the function, f(x), that models Peter's total video game time and explain what each number in the situation represents, using complete sentences.

OpenStudy (anonymous):

@rsadhvika

OpenStudy (anonymous):

not sure how to make an f(x) function lol

OpenStudy (rsadhvika):

Since he playes 5 hours each week, \(\large f(x) = 5x\)

OpenStudy (anonymous):

Alright, and now he's only aloud 5 hours for 8 weeks.

OpenStudy (rsadhvika):

Since he already played 10 hours to start wid, add +10 to above

OpenStudy (rsadhvika):

\(f(x) = 5x + 10\)

OpenStudy (anonymous):

5x + 10

OpenStudy (anonymous):

what about the 8 week limit?

OpenStudy (rsadhvika):

we dont care, that wont come in the equation :) may be use the equation oly for 8 weeks, and after throw it outside the window lol

OpenStudy (anonymous):

lol

OpenStudy (rsadhvika):

\(f(x) = 5x + 10\) , \(x \le \) 8 that tells to use the equation oly for 8 weeks

OpenStudy (anonymous):

wait maybe 8 gets inserted to f(8)?

OpenStudy (rsadhvika):

fine ? :)

OpenStudy (rsadhvika):

nope, nothing like that, 8 is just a constraint for function here

OpenStudy (rsadhvika):

since we dont knw how many hours he is allowed to play after 8 weeks, we say that, the function is GOOD oly for x <= 8

OpenStudy (anonymous):

lol so should i just write f(x) = 5x + 10?

OpenStudy (rsadhvika):

also write the constraint/condition, \(x \le 8\)

OpenStudy (anonymous):

alight, thanks man

OpenStudy (rsadhvika):

the required function for play time is, \( f(x) = 5x + 10 , ~~~ x \le 8\) where x represents the # of weeks passed 10 represents the # of hours he played to start with, that is when x = 0 f(x) gives the # of hours played given the week number x

OpenStudy (rsadhvika):

np :)